OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $6.10. If the number of dimes was tripled and the number of nickels was increased by 49, the value of the coins would be $18.95. How many nickels and dimes does he have?

hero (hero):

.05n + .10d = 6.10 .05(n + 49) + .10(3d) = 18.95 Solve for n and d

OpenStudy (anonymous):

.05N + .10D = 6.10 .05(N+49) + .10*3*D = 18.95 .05(N+49) + .30D = 18.95 -3[.05N + .10D = 6.10] .05N + 2.45 + .30D = 18.95 -.15N - .30D = -18.30 -.10N + 2.45 = .65 2.45 - .65 = .10N 1.80 = .10N N = 1.8/.1 = 18 0.90 in nickels 6.1-.9 = 5.20 52 dimes and 18 nickels

OpenStudy (anonymous):

thats what i got